# DoS Attacks and Defense Technologies in Blockchain Systems: A Hierarchical Analysis

### v1

@@ 1 Introduction
Blockchain technology has made remarkable progress in the international arena and is widely used in other fields such as smart contracts, supply chain, finance, healthcare, and government [10]. This trend of wide application has made it a key technology support in many fields. For example, in the field of smart contracts, blockchain technology brings increased efficiency, transparency, and security. Within the domain of supply chains, it has been shown to promote transparency and improve traceability. Moreover, in the field of finance, blockchain has been found to improve cross-border payments, asset management, and transaction settlement. Finally, in the area of healthcare, it provides a secure framework for managing and sharing health data [20].
@@
(1) Bitcoin Network Transaction Fog Attack. In 2014, the attacker generated a substantial number of small transactions, each comprising numerous inputs and outputs [33]. This led to the accumulation of a large amount of transaction data. These substantial transactions effectively occupied the available space within the block, thus impeding the processing of routine transactions. This phenomenon is referred to as transaction fog, which occurs when the network becomes opaque, hindering the visibility of the actual transactions.
@@
The advent of blockchain technology and the subsequent revelation of its vulnerabilities have prompted the academic community to undertake comprehensive research endeavors concerning its security [32]. In 2018, Cai Z, Du C, and Gan Y provided a concise overview of classification, architecture, and key blockchain technologies. They conducted a preliminary analysis of key management, access control mechanisms, defense mechanisms against DDoS attacks, fragmentation leak prevention mechanisms, and future blockchain development trends. In addition, they predicted the security mechanisms of the blockchain. This analysis was intended to facilitate further research on the blockchain and its security [5]. Subsequently, Gu Xin and Xu Shuzhen et al. expounded on the archetypal blockchain technology architecture, meticulously analyzing the security of blockchain technology from the perspectives of block data structure, hash algorithms, digital signatures, and smart contracts. In subsequent years, a number of researchers and scholars have initiated the discussion of issues pertaining to the security and privacy of Bitcoin and the blockchain [31]. In 2021, Tian et al. conducted a comprehensive analysis of the security and privacy of Bitcoin. Their analysis encompassed various aspects, including hierarchical classification and attack attribution analysis, among others. This study aimed to analyze the system architecture, attack principles, and defense strategies of existing security problems on blockchain by examining the dimensions of Layer Classification and Attack Correlation Analysis [26]. In 2022, Yang et al. presented a comprehensive analysis of 14 prevalent security vulnerabilities in smart contracts and summarized methods of vulnerability-related security prevention. They analyzed four aspects of data storage security, data privacy security, data access security, and data sharing security, which provided some reference for future research work of related people in the field of blockchain security [12].
@@ 2.1.1 Block Structure
Block. The blockchain is composed of blocks as the basic unit and contains metadata. Metadata is extremely important data in a block, including timestamps, hashes of previous blocks, random numbers (Nonce), etc. Each block usually contains data on one or more transactions, which can be digital currency transfers, smart contract executions, etc. The block index is shown in Figure 1.
@@
Hash. The hash algorithm is irreversible, thereby excluding the possibility of recovering the original data from the encrypted version. According to its underlying principle, it can be seen that only slightly changing the data in the block will lead to changes in the hash value. Therefore, the transaction data in the block is encrypted and verified by hashing, which ensures its integrity and security.
Chain. The blockchain is composed of a series of sequentially linked blocks. Each block contains data, including the hash value of the previous block. This results in the data forming a continuous chain on the blockchain. Each node has a complete copy of the blockchain. The blockchain structure from block to chain is shown in Figure 2.
@@
Consensus Mechanism. The consensus mechanisms ensure that all nodes in the network agree to a specific action, thus reaching a consensus. These mechanisms also determine the validity of specific branches, including Proof of Work (PoW), Proof of Stake (PoS), and Proof of Equity Shares (PoA).
@@ 2.1.2 Blockchain Architecture
Cryptography Layer. The cryptography layer, situated at the foundation of the blockchain, employs a range of cryptographic theories, including encryption algorithms, digital signatures, and hash functions, to ensure the security and confidentiality of data and communications within the blockchain system. The implementation of encryption technology ensures the confidentiality, integrity, and protection of transaction data, protecting against potential threats such as tampering and theft.
@@
Data Layer. The data layer is widely involved in the database, transaction data, smart contract code, and other important information in the blockchain. It is responsible for the management of data storage and retrieval in the blockchain system. The data layer also includes the processing of encryption, compression, and indexing of data, with data storage and retrieval functions to ensure the security and integrity of data.
@@
Network Layer. The network layer encompasses the network topology, routing protocols, peer-to-peer communication protocols, and other related components. The primary function is to facilitate the establishment and maintenance of connections and communications between nodes. This plays a pivotal role in determining the scalability, security, and performance of the blockchain network. The system must provide a secure and reliable communication environment to ensure that data can be exchanged and consensus reached between nodes.
@@
Consensus Layer. The consensus layer is the core content of the blockchain system, including the consensus algorithm, the rules of the block generation, and the verification mechanism, providing the technical infrastructure for the blockchain. The design of the consensus layer affects the security, performance, and decentralization of the blockchain system. It defines the rules and algorithms for reaching consensus among nodes and ensures that all nodes agree on the state of the blockchain.
@@
Incentive Layer. The incentive layer includes incentives and reward rules that are designed to facilitate desired behaviors. Common incentive mechanisms include block rewards and transaction fees. Through the design of the incentive layer, it can promote the behavior of nodes in line with the interests of the system and guarantee the security and stability of the blockchain network.
@@
Contract Layer. The contract layer includes the execution environment and related tools for smart contracts. It is mainly responsible for processing and executing the code of smart contracts. The existence of the contract layer provides a secure and trustworthy execution environment and ensures the correct execution and non-tamperability of smart contracts.
@@
Application Layer. The application layer is the interface for users to interact with the blockchain system, including the user interface, applications, and smart contracts. Users can interact with the blockchain system through applications, sending transactions, querying data, and other operations. The application layer provides users with a way to access the blockchain network, enabling blockchain technology to be applied in various scenarios, such as cryptocurrency trading and supply chain management.
@@ 2.2 Common Security Threats
Double Spending. Double spending is defined as the transmission of the same digital currency to two distinct addresses in a single transaction. This occurrence is typically attributable to network delays or exploitation of system vulnerabilities by attackers. This phenomenon may lead to a reduction in system trust, thereby eroding user confidence and, consequently, exerting an adverse effect on the overall reliability and trustworthiness of the blockchain system.
51% Attack. 51% attack is characterized by the attacker’s possession of more than half of the computing power in the network, thereby enabling them to exert control over the consensus process within the network. The repercussions of these actions can encompass a range of issues, including the manipulation of transactions, the occurrence of double spending, and the disruption of service. In particularly grave instances, the integrity and stability of the entire blockchain network may be compromised.
Smart Contract Vulnerability. Smart contracts are based on code execution, and the existence of coding errors or vulnerabilities may lead to asset theft, contract deadlock, or abnormal execution. The ramifications of these actions include the forfeiture of assets, the execution of contracts that deviate from standard practices, and the compromise of system security. These factors exert a direct influence on the stability of the blockchain system and the confidence of its users.
Loss or leakage of the private key. The private key is how the digital asset is controlled. If the private key is lost or disclosed, attackers may gain control of the asset, which may result in theft of the asset, impersonation of the user’s identity, and loss of control of the asset by the legitimate user.
Leakage of transaction information. Transaction information on blockchain is usually public, but sometimes users want to protect their privacy. After the transaction information is leaked, attackers can track the user’s identity and behavior through their analysis of the transaction information, which leads to the exposure of the user’s privacy, identity theft, and tracking of transaction behavior, and then violates the user’s rights and interests.
Fork Attack. Fork attack refers to the occurrence of a fork in the blockchain network, which is used by attackers to implement malicious behaviors such as double spending and canceling transactions. It may lead to chaotic transactions, compromised asset security, and reduced system credibility, directly threatening the stability and reliability of the blockchain system.
Malicious Smart Contracts. Malicious smart contracts are contracts designed to attack users, which may include theft of assets, denial of service, market manipulation, and other malicious behaviors, which can lead to loss of assets, damage to users’ rights and interests, and decline in system trust, directly affecting the security of the blockchain system and the interests of users.
@@ 2.3 DoS Attack
DoS attack is a relatively low-cost form of cyberattack with the potential to have more severe consequences. Attackers can use relatively simple tools and techniques to cause network performance degradation, node unavailability, smart contract unavailability, and data loss in blockchain systems.
@@ 2.4.1 SYN Flood Attack
SYN flood attack is a common DoS attack, in which the attacker exploits a vulnerability in the TCP protocol’s three-handshake process by sending a large number of forged SYN packets to the target server, but does not send subsequent ACK packets. Since the server receives a large number of connection requests but is unable to complete the three handshakes, it will remain connected while waiting for the client’s ACK packet. The attacker keeps sending forged connection requests, which makes a large number of half-connected states (SYN_RCVD states) appear on the server, consuming server resources and ultimately causing the server to be unable to respond to normal user requests [11].
@@
  • Limit the number of connections. Set firewall rules or server parameters to limit the number of connections to a single IP address to prevent too many connections from a single IP.
@@ 2.4.2 ICMP Attack
The ICMP protocol is used to send diagnostic and error messages, while ICMP attacks utilize some features or vulnerabilities of the ICMP protocol to attack the target system, including destination unreachable, timeout, etc. The attacker can use these messages to send a large number of false ICMP messages, thus consuming the network bandwidth and processing power of the target system and causing network congestion or resource exhaustion of the target system [19].
@@ 2.4.3 Reflection Amplification Attack
  • Update and configure servers. Update and configure servers regularly to patch known vulnerabilities and weaknesses and reduce the attack surface.
@@ 3 DoS Attacks in Blockchain Systems
There are two reasons why traditional DoS attacks cannot be employed against blockchain networks. Firstly, nodes are dynamic and can access new nodes at any time, making it difficult for an attacker to find all nodes. Secondly, the presence of tens of thousands of nodes complicates the ability of an attacker to launch an attack on all of them simultaneously. Consequently, achieving a complete denial of service in blockchain systems is difficult. Instead, the objective of DoS attacks is to compromise or disrupt the functionality of the blockchain system.
@@ 3.1 Network Layer
Denial-of-service attacks at the blockchain network layer aim to prevent the proper functioning of the blockchain network by taking up network bandwidth, disrupting communication between nodes, or rendering nodes inoperable, preventing communication between nodes and causing network fragmentation or delays.
@@ 3.1.2 Distributed Denial of Service
Denial-of-service attacks typically necessitate the utilization of a substantial amount of network resources by the attacker to attack the target network or computer system. This results in the target system’s inability to respond promptly to normal access, or even the system’s failure. However, in the context of the blockchain system, the user nodes are tens of thousands, and the resources of each node are limited. Consequently, the attacker generally opts to integrate network bandwidth fragmentation to implement distributed denial-of-service (DDoS) attacks. The principle of DDoS attacks based on the blockchain network is shown in Figure 4.
@@ 3.2 Consensus Layer
Denial-of-service attacks at the blockchain consensus layer aim to disrupt or interfere with the consensus mechanism of the blockchain network, thereby preventing nodes from agreeing on the correct blockchain state. The most prevalent attack methods include the 51% attack and the network splitting attack. In this section, we methodically analyze the 51% attack as a paradigmatic case.

### v2

@@ 1 Introduction
Blockchain technology has made remarkable progress in the international arena and is widely used in other fields such as smart contracts, supply chain, finance, healthcare, and government [18]. This trend of wide application has made it a key technology support in many fields. For example, in the field of smart contracts, blockchain technology brings increased efficiency, transparency, and security. Within the domain of supply chains, it has been shown to promote transparency and improve traceability [5]. Moreover, in the field of finance, blockchain has been found to improve cross-border payments, asset management, and transaction settlement. Finally, in the area of healthcare, it provides a secure framework for managing and sharing health data [34].
@@
(1) Bitcoin Network Transaction Fog Attack. In 2014, the attacker generated a substantial number of small transactions, each comprising numerous inputs and outputs [59]. This led to the accumulation of a large amount of transaction data. These substantial transactions effectively occupied the available space within the block, thus impeding the processing of routine transactions. This phenomenon is referred to as transaction fog, which occurs when the network becomes opaque, hindering the visibility of the actual transactions [58].
@@
The advent of blockchain technology and the subsequent revelation of its vulnerabilities have prompted the academic community to undertake comprehensive research endeavors concerning its security [57]. In 2018, Cai Z, Du C, and Gan Y provided a concise overview of classification, architecture, and key blockchain technologies [13]. They conducted a preliminary analysis of key management, access control mechanisms, defense mechanisms against DDoS attacks, fragmentation leak prevention mechanisms, and future blockchain development trends. In addition, they predicted the security mechanisms of the blockchain. This analysis was intended to facilitate further research on the blockchain and its security [8]. Subsequently, Gu Xin and Xu Shuzhen et al. expounded on the archetypal blockchain technology architecture, meticulously analyzing the security of blockchain technology from the perspectives of block data structure, hash algorithms, digital signatures, and smart contracts. In subsequent years, a number of researchers and scholars have initiated the discussion of issues pertaining to the security and privacy of Bitcoin and the blockchain [56]. In 2021, Tian et al. conducted a comprehensive analysis of the security and privacy of Bitcoin. Their analysis encompassed various aspects, including hierarchical classification and attack attribution analysis, among others. This study aimed to analyze the system architecture, attack principles, and defense strategies of existing security problems on blockchain by examining the dimensions of Layer Classification and Attack Correlation Analysis [44]. In 2022, Yang et al. presented a comprehensive analysis of 14 prevalent security vulnerabilities in smart contracts and summarized methods of vulnerability-related security prevention [47]. They analyzed four aspects of data storage security, data privacy security, data access security, and data sharing security, which provided some reference for future research work of related people in the field of blockchain security [20].
@@ 2.1.1 Block Structure
Block. The blockchain is composed of blocks as the basic unit and contains metadata. Metadata is extremely important data in a block, including timestamps, hashes of previous blocks, random numbers (Nonce), etc [9]. Each block usually contains data on one or more transactions, which can be digital currency transfers, smart contract executions, etc. The block index is shown in Figure 1.
@@
Hash. The hash algorithm is irreversible, thereby excluding the possibility of recovering the original data from the encrypted version [24]. According to its underlying principle, it can be seen that only slightly changing the data in the block will lead to changes in the hash value. Therefore, the transaction data in the block is encrypted and verified by hashing, which ensures its integrity and security.
Chain. The blockchain is composed of a series of sequentially linked blocks. Each block contains data, including the hash value of the previous block [1]. This results in the data forming a continuous chain on the blockchain. Each node has a complete copy of the blockchain. The blockchain structure from block to chain is shown in Figure 2.
@@
Consensus Mechanism. The consensus mechanisms ensure that all nodes in the network agree to a specific action, thus reaching a consensus [45]. These mechanisms also determine the validity of specific branches, including Proof of Work (PoW), Proof of Stake (PoS), and Proof of Equity Shares (PoA).
@@ 2.1.2 Blockchain Architecture
Cryptography Layer. The cryptography layer, situated at the foundation of the blockchain, employs a range of cryptographic theories, including encryption algorithms, digital signatures, and hash functions, to ensure the security and confidentiality of data and communications within the blockchain system [50]. The implementation of encryption technology ensures the confidentiality, integrity, and protection of transaction data, protecting against potential threats such as tampering and theft.
@@
Data Layer. The data layer is widely involved in the database, transaction data, smart contract code, and other important information in the blockchain. It is responsible for the management of data storage and retrieval in the blockchain system. The data layer also includes the processing of encryption, compression, and indexing of data, with data storage and retrieval functions to ensure the security and integrity of data [38].
@@
Network Layer. The network layer encompasses the network topology, routing protocols, peer-to-peer communication protocols, and other related components. The primary function is to facilitate the establishment and maintenance of connections and communications between nodes [25]. This plays a pivotal role in determining the scalability, security, and performance of the blockchain network. The system must provide a secure and reliable communication environment to ensure that data can be exchanged and consensus reached between nodes.
@@
Consensus Layer. The consensus layer is the core content of the blockchain system, including the consensus algorithm, the rules of the block generation, and the verification mechanism, providing the technical infrastructure for the blockchain [14]. The design of the consensus layer affects the security, performance, and decentralization of the blockchain system. It defines the rules and algorithms for reaching consensus among nodes and ensures that all nodes agree on the state of the blockchain.
@@
Incentive Layer. The incentive layer includes incentives and reward rules that are designed to facilitate desired behaviors. Common incentive mechanisms include block rewards and transaction fees. Through the design of the incentive layer, it can promote the behavior of nodes in line with the interests of the system and guarantee the security and stability of the blockchain network [17].
@@
Contract Layer. The contract layer includes the execution environment and related tools for smart contracts. It is mainly responsible for processing and executing the code of smart contracts. The existence of the contract layer provides a secure and trustworthy execution environment and ensures the correct execution and non-tamperability of smart contracts [35].
@@
Application Layer. The application layer is the interface for users to interact with the blockchain system, including the user interface, applications, and smart contracts. Users can interact with the blockchain system through applications, sending transactions, querying data, and other operations. The application layer provides users with a way to access the blockchain network, enabling blockchain technology to be applied in various scenarios, such as cryptocurrency trading and supply chain management [23].
@@ 2.2 Common Security Threats
Double Spending. Double spending is defined as the transmission of the same digital currency to two distinct addresses in a single transaction. This occurrence is typically attributable to network delays or exploitation of system vulnerabilities by attackers [40]. This phenomenon may lead to a reduction in system trust, thereby eroding user confidence and, consequently, exerting an adverse effect on the overall reliability and trustworthiness of the blockchain system.
51% Attack. 51% attack is characterized by the attacker’s possession of more than half of the computing power in the network, thereby enabling them to exert control over the consensus process within the network. The repercussions of these actions can encompass a range of issues, including the manipulation of transactions, the occurrence of double spending, and the disruption of service. In particularly grave instances, the integrity and stability of the entire blockchain network may be compromised [49].
Smart Contract Vulnerability. Smart contracts are based on code execution, and the existence of coding errors or vulnerabilities may lead to asset theft, contract deadlock, or abnormal execution. The ramifications of these actions include the forfeiture of assets, the execution of contracts that deviate from standard practices, and the compromise of system security. These factors exert a direct influence on the stability of the blockchain system and the confidence of its users [21].
Loss or leakage of the private key. The private key is how the digital asset is controlled. If the private key is lost or disclosed, attackers may gain control of the asset, which may result in theft of the asset, impersonation of the user’s identity, and loss of control of the asset by the legitimate user [43].
Leakage of transaction information. Transaction information on blockchain is usually public, but sometimes users want to protect their privacy. After the transaction information is leaked, attackers can track the user’s identity and behavior through their analysis of the transaction information, which leads to the exposure of the user’s privacy, identity theft, and tracking of transaction behavior, and then violates the user’s rights and interests [49].
Fork Attack. Fork attack refers to the occurrence of a fork in the blockchain network, which is used by attackers to implement malicious behaviors such as double-spending and canceling transactions. It may lead to chaotic transactions, compromised asset security, and reduced system credibility, directly threatening the stability and reliability of the blockchain system [40].
Malicious Smart Contracts. Malicious smart contracts are contracts designed to attack users, which may include theft of assets, denial of service, market manipulation, and other malicious behaviors, which can lead to loss of assets, damage to users’ rights and interests, and decline in system trust, directly affecting the security of the blockchain system and the interests of users [23].
@@ 2.3 DoS Attack
DoS attack is a relatively low-cost form of cyberattack with the potential to have more severe consequences [16]. Attackers can use relatively simple tools and techniques to cause network performance degradation, node unavailability, smart contract unavailability, and data loss in blockchain systems.
@@ 2.4.1 SYN Flood Attack
SYN flood attack is a common DoS attack, in which the attacker exploits a vulnerability in the TCP protocol’s three-handshake process by sending a large number of forged SYN packets to the target server, but does not send subsequent ACK packets [22]. Since the server receives a large number of connection requests but is unable to complete the three handshakes, it will remain connected while waiting for the client’s ACK packet. The attacker keeps sending forged connection requests, which makes a large number of half-connected states (SYN_RCVD states) appear on the server, consuming server resources and ultimately causing the server to be unable to respond to normal user requests [19].
@@
  • Limit the number of connections. Set firewall rules or server parameters to limit the number of connections to a single IP address to prevent too many connections from a single IP [54].
@@ 2.4.2 ICMP Attack
The ICMP protocol is used to send diagnostic and error messages, while ICMP attacks utilize some features or vulnerabilities of the ICMP protocol to attack the target system, including destination unreachable, timeout, etc [2]. The attacker can use these messages to send a large number of false ICMP messages, thus consuming the network bandwidth and processing power of the target system and causing network congestion or resource exhaustion of the target system [33].
@@ 2.4.3 Reflection Amplification Attack
  • Update and configure servers. Update and configure servers regularly to patch known vulnerabilities and weaknesses and reduce the attack surface [15].
@@ 3 DoS Attacks in Blockchain Systems
There are two reasons why traditional DoS attacks cannot be employed against blockchain networks [27]. Firstly, nodes are dynamic and can access new nodes at any time, making it difficult for an attacker to find all nodes. Secondly, the presence of tens of thousands of nodes complicates the ability of an attacker to launch an attack on all of them simultaneously. Consequently, achieving a complete denial of service in blockchain systems is difficult. Instead, the objective of DoS attacks is to compromise or disrupt the functionality of the blockchain system.
@@ 3.1 Network Layer
Denial-of-service attacks at the blockchain network layer aim to prevent the proper functioning of the blockchain network by taking up network bandwidth, disrupting communication between nodes, or rendering nodes inoperable, preventing communication between nodes and causing network fragmentation or delays [51].
@@ 3.1.2 Distributed Denial of Service
Denial-of-service attacks typically necessitate the utilization of a substantial amount of network resources by the attacker to attack the target network or computer system. This results in the target system’s inability to respond promptly to normal access, or even the system’s failure [46]. However, in the context of the blockchain system, the user nodes are tens of thousands, and the resources of each node are limited. Consequently, the attacker generally opts to integrate network bandwidth fragmentation to implement distributed denial-of-service (DDoS) attacks. The principle of DDoS attacks based on the blockchain network is shown in Figure 4.
@@ 3.2 Consensus Layer
Denial-of-service attacks at the blockchain consensus layer aim to disrupt or interfere with the consensus mechanism of the blockchain network, thereby preventing nodes from agreeing on the correct blockchain state [10]. The most prevalent attack methods include the 51% attack and the network splitting attack. In this section, we methodically analyze the 51% attack as a paradigmatic case.